# Latent Gaussian dynamic factor modeling and forecasting for multivariate count time series

### v2

@@ 3.2 Estimating link function and calculating its inverse
so that one has the points $(v_{m},L_{ij}^{-1}(v_{m}))=(v_{m},u_{m})$ on the curve $L_{ij}^{-1}(v)=u$. The value of $L_{ij}^{-1}(v)$ for other points $v$ can then be obtained through some interpolation, for example, natural cubic splines (e.g. Chapter 8 in Kress, , 1998). The natural cubic splines produce piecewise polynomial functions, having smooth derivatives. In addition, we use finer grids for $u$ near ±$\pm 1$ while wider grids are used around $u=0$.
From a numerical standpoint, the $M+1$ points $(v_{m},u_{m})$ satisfy $v_{m}=L_{ij}^{-1}(u_{m})$, and the interpolation of $L_{ij}^{-1}$ is defined as
$\widetilde{L}_{ij}^{-1}(v)=:\sum_{m=1}^{M}\widetilde{L}_{ij,m}^{-1}(v)\bm{1}_{[v_{m-1},v_{m})}(v)$
for $v\in(v_{m},v_{m+1})$, $m=0,1,\ldots,M$, with $M$ pieces of the 3rd-order spline polynomials
$\widetilde{L}_{ij,m}^{-1}(v)=a_{m-1}\frac{(v_{m}-v)^{3}}{6h_{m}}+a_{m}\frac{(v-v_{m-1})^{3}}{6h_{m}}+b_{1,m}(v-v_{m-1})+b_{2,m}(v_{m}-v),$ (3.13)
where $h_{m}=v_{m}-v_{m-1}$. The polynomials in (3.13) satisfy the following constraints,
$\widetilde{L}_{ij,m}^{-1}(v_{m-1})=u_{m-1},\quad\widetilde{L}_{ij,m}^{-1}(v_{m})=u_{m},\quad m=1,\ldots,M,$ (3.14)
$(\widetilde{L}_{ij,m}^{-1})^{\prime}(v_{m})=(\widetilde{L}_{ij,m+1}^{-1})^{\prime}(v_{m}),\quad m=1,\ldots,M-1,$ (3.15)
$(\widetilde{L}_{ij,m}^{-1})^{\prime\prime}(v_{m})=(\widetilde{L}_{ij,m+1}^{-1})^{\prime\prime}(v_{m}),\quad m=1,\ldots,M-1,$ (3.16)
$(\widetilde{L}_{ij,1}^{-1})^{\prime\prime}(-v_{0})=(\widetilde{L}_{ij,n}^{-1})^{\prime\prime}(v_{M})=0.$ (3.17)
Note that the knots in (3.14) form a continuity of the spline function, (3.15) and (3.16) ensure the second-order smoothness. (3.17), called the natural boundary, is required for determining polynomials uniquely. We extend the interpolation further to $[-1,1]$ by letting $\widetilde{L}_{ij}^{-1}(v)=-1$ for $v\in[-1,v_{0})$ and $\widetilde{L}_{ij}^{-1}(v)=1$ for $v\in(v_{M},1]$. Since the equations (3.13) can be written as a symmetric tridiagonal system, it can be easily solved.
Figure 1 depicts $L_{ij}(u),L_{ij}^{-1}(v)$ and its interpolation $\widetilde{L}_{ij}^{-1}(v)$ for four representative marginal count distributions: Bernoulli, multinomial, Poisson, and negative binomial. For example, the Bernoulli case considers a pair of $F_{i}=\textrm{Bern}(p_{i})$ and $F_{j}=\textrm{Bern}(p_{j})$ with four different choices of combinations $(p_{i},p_{j})$. Several combinations of parameters for other types of distributions, including multinomial, Poisson, and negative binomial distributions, are also considered. As seen from the plots, the inverse link function $L_{ij}^{-1}$ obtained by flipping the axes and numerical inverse $\widetilde{L}_{ij}^{-1}$ obtained through the interpolation are nearly indistinguishable. These combinations of marginal distributions are used in the simulation study in Section 6.
3.3 Selecting the number of factors and lag order of factor series
In practice, the number of factor series $r$ and their lag order $p$ in (2.4) are unknown and need to be chosen for the model estimation. We consider here several methods for this task based on available approaches and also introduced cross-validation schemes tailored for our model. Their performance is assessed in Section 6.1.2 below, with cross-validation schemes generally outperforming others.
3.3.1 Selection of $r$
@@
3.3.2 Selection of $p$
One common approach is to use an information criterion,
$\widehat{p}=\mathop{\rm argmin}_{l}\left\{\ln(|\widehat{\Sigma}_{\eta}(l)|)+g_{i,l}(r,T)\right\},$ (3.21)
$\begin{gathered}g_{1,l}(r,T)=\frac{2}{T}lr^{2},\quad g_{2,l}(r,T)=\frac{2\log(\log(T))}{T}lr^{2},\\ g_{3,l}(r,T)=\frac{\log(T)}{T}lr^{2},\quad g_{4,l}(r,T)=\frac{2r(rl+1)}{T}.\end{gathered}$ (3.22)
The resulting criteria can be found in Chapter 4.3 of Lütkepohl, (2005).
We now propose a cross-validation strategy tailored to our model. We shall exploit once again the idea of Remark 3.1. Were the factor series $Y_{t}$ observed, a natural cross-validation scheme would select $p$ as $l$ minimizing
the function to minimize can be replaced by
or
$\sum_{b=1}^{B}\left(-2\widehat{\bm{\psi}}^{(-b,l)^{\prime}}\widehat{\gamma}_{Y}^{(b,l)}+\widehat{\bm{\psi}}^{(-b,l)^{\prime}}\widehat{\Gamma}_{Y}^{(b,l)}\widehat{\bm{\psi}}^{(-b,l)}\right),$ (3.23)
$\widehat{\gamma}_{Y}^{(b,l)}=\operatorname{vec}(\mathcal{X}^{(b,l)^{\prime}}\mathcal{Y}^{(b,l)}),\quad\widehat{\Gamma}_{Y}^{(b,l)}=I_{r}\otimes\mathcal{X}^{(b,l)^{\prime}}\mathcal{X}^{(b,l)}.$ (3.24)
Following the idea of Remark 3.1, note that the quantities in (3.24) are obtained from the (sample) second-order properties of $(Y_{t}^{(b)})$. For our model, they can be replaced by those implied by the equation (2.12) based on the observed data $(X_{t}^{(b)})$. Similarly, we already have the estimators $\widehat{\Psi}^{(-b,l)}$ calculated from the observed data $(X_{t}^{(-b)})$. In summary, as a cross-validation scheme to select $p$ for our model, we also minimize (3.23) but where the quantities involved are computed on the training data $(X_{t}^{(-b)})$ and the testing data $(X_{t}^{(b)})$.
Remark 3.2.
Note that determining the number of factors can be performed regardless of the lag order of the factor series. We thus recommend choosing $r$ first, followed by selecting $p$.
@@ 4 Theoretical properties
In this section, we investigate theoretical properties of our estimators for the latent factor model (2.3). Our results are derived for the PCA estimators of the latent factor model in Section 3. The proofs are based on Düker et al., (2023) who provide a theoretical foundation to investigate the model (2.1) for general latent Gaussian processes. We further base our investigations on results in Doz et al., (2011) who prove that the PCA estimators are consistent when the series following a factor model is observed and also on concentration results of bounded functions of Markov chains in Fan et al., (2021). We start with briefly recalling our estimation procedure and introduce some additional notation in Section 4.1. We collect assumptions in Section 4.2 and then state our main results in Section 4.3. Section 4.4 concludes with a particular case satisfying the assumptions made in Sections 4.2.
4.1 Estimation
@@ 4.2 Assumptions
Assumption L.3 is shown to hold for several common count distributions under Assumption L.2; see Appendix E in Düker et al., (2023). Note that we require our moment conditions to hold uniformly in a neighborhood around $\theta_{i}$. This allows us to infer finiteness on a compact subset of the parameter space of $\theta_{i}$. The second set of assumptions is needed to derive consistency results for estimators of the latent factor model.
@@
We remark also that Assumptions C.1 and C.2 are the analogues of Assumptions C.1 and C.2 in Düker et al., (2023). Assumptions C.1 and C.2 in Düker et al., (2023) ensure consistent estimation of $R_{Z}$ when $\{Z_{t}\}$ follows a $d-$dimensional VAR($p$) model. Those assumptions are concentration inequalities, suggesting non-asymptotic bounds for the deviation between estimated and true quantities. We use here a representation in terms of stochastic boundedness which is a weaker assumption but sufficient for our purposes since we aim to recover the results in Doz et al., (2011) who also state their results in terms of stochastic boundedness.
@@
$\{Y_{t}\}$ and $\{\varepsilon_{t}\}$ are independent and:
  • $r-$vector factor series $Y_{t}=(Y_{i,t})_{i=1,\ldots,r}$ follows a stationary VAR($p$) model given by
@@ 5.1 Particle-based sampling procedure
The following presentation extends that of Jia et al., (2023) to the multivariate setting relevant to the Monte Carlo approximation problem. The quantity (5.3) is known to be well approximated through Sequential Monte Carlo (SMC) by generating “particles” over time $t$ (e.g. Doucet et al., , 2001, Doucet and Johansen, , 2009). The main difference from Jia et al., (2023) is that the model here has two latent processes $\{Z_{t}\}$ and $\{Y_{t}\}$. To deal with the two latent processes, we use Kalman recursions to forecast and update the latent process $\{Y_{t}\}$ and approximate the distribution of $\{Z_{t}\}$ conditioning on $\{X_{t}\}$. This approach is called Rao-Blackwellization and the method is adapted from the partially observed Gaussian state-space models (e.g. Andrieu and Doucet, , 2002, Briers et al., , 2010).
The following is the SMC algorithm for particle filtering to generate particles $\{\widetilde{Z}_{t}^{(k)}\}_{t=1,\ldots,T}$, $k=1,\ldots,N$, over time, whose weighted average then approximates (5.3). The particle $\{\widetilde{Z}_{t}^{(k)}\}$ can be regarded as the realization of the underlying latent process. Additionally, we add a resampling step which is often used in sequential Monte Carlo algorithms. We will explain the necessity of resampling below.
@@
    $w_{t}(\widehat{Z}_{t|t-1}^{(k)})=\mathbb{P}\left(\mathcal{N}(\widehat{Z}_{t|t-1}^{(k)},\widehat{R}_{t|t-1})\in A_{x_{t}}\right).$
@@
with $\sum_{j=1}^{0}=0$ as a convention. This is used in Step 3 of the SIS/R algorithm above. Alternatively, one can use multinomial resampling, which is resampling particles by regarding $\{\widetilde{w}_{t}^{(k)}\}$ as a multinomial probability distribution. Many other resampling methods exist (see Douc and Cappé, (2005) for more information).
@@ 5.3 On forecasting for longer horizons
In this section, we briefly discuss what to expect from the forecasting method when the forecasting horizon $h$ becomes longer. Recall that the latent factor series is stationary and follows a stable VAR model. The latent process $\{Z_{t}\}$ is also stationary and its long-term prediction converges to its mean, which is a zero vector. From (B.13) and (B.15), the predicted particles are therefore expected to converge eventually to zero vectors as well. Thus, for longer horizon $h$, we expect
@@
First, in some instances, e.g. Bernoulli and multinomial distributions, $\theta_{i}$ represents proportion of count values and is estimated as the corresponding sample proportions. In these cases, for long horizon $h$, we therefore expect our forecast to yield the most likely observed count (modulo the issue of ties). This is the case with the application considered in Section 7. On the other hand, for many other distributions, this observation may not necessarily hold. For example, for the Poisson distribution, the parameter is taken as the sample mean and the most likely count according to this Poisson distribution does not need to be the most likely observed value, let alone be in the sample.
@@ 6 Simulation study
In this section, we assess the performance of estimation and forecasting procedures introduced in Sections 3 and 5. We focus on Bernoulli, multinomial, Poisson, and negative binomial marginal distributions with several parameter values. Our goal is to check how the performance of our proposed methods compares to that of the traditional PCA methods in the factor modeling literature (e.g., Lam et al., , 2011, Doz et al., , 2011, 2012).
@@ 6.1.1 Estimation for known numbers of factors and lag order
The simulation settings are as follows. First, we fix the order $p=1$ for factor series $\{Y_{t}\}$ in (2.4) and take the number of factor series as $r=2$ or $r=5$. These values are assumed to be given in this section. Then, by following the standard DFM literature (e.g., Doz et al., , 2011), we generate the latent Gaussian DFM (2.3)–(2.4) with parameters:
@@
  • For multinomial marginal distributions with $\theta_{i}=(p_{1,i},\ldots,p_{5,i})$ where the number of the categories is given as 5, $\theta_{i}=(0.2,\ldots,0.2)$ (uniform), $(0,0.25,0.5,0.25,0)$ (unimodal), and $(0.45,0,0.1,0,0.45)$ (trimodal) for $i\in I_{1}$, $I_{2}$, and $I_{3}$, respectively.
@@
6.1.2 Selection of the number of factors and lag order

### v3

@@ 3.2 Calculating inverse of link function
so that one has the points $(v_{m},L_{ij}^{-1}(v_{m}))=(v_{m},u_{m})$ on the curve $L_{ij}^{-1}(v)=u$. The value of $L_{ij}^{-1}(v)$ for other points $v$ can then be obtained through piecewise linear interpolation. In addition, we use finer grids for $u$ near ±$\pm 1$ while wider grids are used around $u=0$. From a numerical standpoint, the $M+1$ points $(v_{m},u_{m})$ satisfy $v_{m}=L_{ij}^{-1}(u_{m})$, and the interpolation of $L_{ij}^{-1}$ is defined as
$\widetilde{L}_{ij}^{-1}(v):=\sum_{m=1}^{M}\widetilde{L}_{ij,m}^{-1}(v)\bm{1}_{[v_{m-1},v_{m})}(v)$
for $v\in(v_{m},v_{m+1})$, $m=0,1,\ldots,M$, with $M$ pieces of the linear spline polynomials
$\widetilde{L}_{ij,m}^{-1}(v)=\widetilde{L}_{ij,m}^{-1}(v_{m-1})+\frac{\widetilde{L}_{ij,m}^{-1}(v_{m})-\widetilde{L}_{ij,m}^{-1}(v_{m-1})}{v_{m}-v_{m-1}}(v-v_{m-1}).$
Figure 1 depicts $L_{ij}(u),L_{ij}^{-1}(v)$ and its interpolation $\widetilde{L}_{ij}^{-1}(v)$ for four representative marginal count distributions: Bernoulli, categorical, Poisson, and negative binomial. For example, the Bernoulli case considers a pair of $F_{i}=\textrm{Bern}(p_{i})$ and $F_{j}=\textrm{Bern}(p_{j})$ with four different choices of combinations $(p_{i},p_{j})$. Several combinations of parameters for other types of distributions, including categorical, Poisson, and negative binomial distributions, are also considered. As seen from the plots, the inverse link function $L_{ij}^{-1}$ obtained by flipping the axes and numerical inverse $\widetilde{L}_{ij}^{-1}$ obtained through the linear interpolation are nearly indistinguishable. These combinations of marginal distributions are used in the simulation study in Section 6.
3.3 Selecting the number of factors
In practice, the number of factor series $r$ is unknown and needs to be chosen for the model estimation. We consider here several methods for this task based on available approaches and also introduce cross-validation schemes tailored for our model.
@@ 3.3 Selecting the number of factors
The selection of lag order $p$ in (2.4) can be similarly conducted by using cross-validation, which is explained in Supplemental Material. Note that determining the number of factors can be performed regardless of the lag order of the factor series. We thus recommend choosing $r$ first, followed by selecting $p$.
@@ 4 Theoretical properties
In this section, we investigate theoretical properties of our estimators for the latent factor model (2.3). Our results are derived for the PCA estimators of the latent factor model in Section 3. The proofs are based on Düker et al., (2024) who provide a theoretical foundation to investigate the model (2.1) for general latent Gaussian processes. We base our investigations on results in Doz et al., (2011) who prove that the PCA estimators are consistent when the series following a factor model is observed and also on concentration results of bounded functions of Markov chains in Fan et al., (2021). We start with briefly recalling our estimation procedure and introduce some additional notation in Section 4.1. We collect assumptions in Section 4.2 and then state our main results in Section 4.3. Section 4.4 concludes with a particular case satisfying the assumptions made in Section 4.2.
4.1 Estimation procedure with additional notation
@@ 4.2 Assumptions
The constant $\bm{c}_{Z}$ in Assumption L.1 controls the dependence of the series $\{Z_{t}\}$ and appears in the non-asymptotic bound derived in Düker et al., (2024) which are used in our work. The assumption is expected to hold for some $\bm{c}_{Z}$ for the processes considered both here and in Düker et al., (2024). Assumption L.3 is shown to hold for several common count distributions under Assumption L.2; see Appendix E in Düker et al., (2024). Note that we require our moment conditions to hold uniformly in a neighborhood around $\theta_{i}$. This allows us to infer finiteness on a compact subset of the parameter space of $\theta_{i}$. The second set of assumptions is needed to derive consistency results for estimators of the latent factor model. The asymptotics described below are understood as $d,T\to\infty$.
@@
We remark also that Assumptions C.1 and C.2 are the analogues of Assumptions C.1 and C.2 in Düker et al., (2024). Assumptions C.1 and C.2 in Düker et al., (2024) ensure consistent estimation of $R_{Z}$ when $\{Z_{t}\}_{t\in\mathbb{Z}}$ follows a $d-$dimensional VAR($p$) model. Those assumptions are concentration inequalities, suggesting non-asymptotic bounds for the deviation between estimated and true quantities. We use here a representation in terms of stochastic boundedness which is a weaker assumption but sufficient for our purposes since we aim to recover the results in Doz et al., (2011) who also state their results in terms of stochastic boundedness.
@@
$\{Y_{t}\}_{t\in\mathbb{Z}}$ and $\{\varepsilon_{t}\}_{t\in\mathbb{Z}}$ are independent and:
  • $r-$vector factor series $\{Y_{t}\}_{t\in\mathbb{Z}}=\{(Y_{i,t})_{i=1,\ldots,r}\}_{t\in\mathbb{Z}}$ follows a stationary VAR($p$) model given by
@@ 5.1 Particle-based sampling procedure
The following presentation extends that of Jia et al., (2023) to the multivariate setting relevant to the Monte Carlo approximation problem. The quantity (5.3) is known to be well approximated through Sequential Monte Carlo (SMC) by generating “particles” over time $t$ (e.g. Doucet et al., , 2001, Doucet and Johansen, , 2009). The main difference from Jia et al., (2023) is that the model here has two latent processes $\{Z_{t}\}_{t\in\mathbb{Z}}$ and $\{Y_{t}\}_{t\in\mathbb{Z}}$. To deal with the two latent processes, we use Kalman recursions to forecast and update the latent process $\{Y_{t}\}$ and approximate the distribution of $\{Z_{t}\}$ conditioning on $\{X_{t}\}$. This approach is called Rao-Blackwellization and the method is adapted from the partially observed Gaussian state-space models (e.g. Andrieu and Doucet, , 2002, Briers et al., , 2010).
The following is the SMC algorithm for particle filtering to generate particles $\{\widetilde{Z}_{t}^{(k)}\}_{t=1,\ldots,T}$, $k=1,\ldots,N$, over time, whose weighted average then approximates (5.3). The particle $\{\widetilde{Z}_{t}^{(k)}\}_{t=1,\ldots,T}$ can be regarded as the realization of the underlying latent process. Additionally, we add a resampling step which is often used in sequential Monte Carlo algorithms. We will explain the necessity of resampling below.
@@
    $w_{t}(\widehat{Z}_{t|t-1}^{(k)})=\mathbb{P}\left(\mathcal{N}(\widehat{Z}_{t|t-1}^{(k)},\widehat{R}_{t|t-1})\in A_{x_{t}}\right).$ (5.6)
@@
with $\sum_{j=1}^{0}=0$ as a convention. This is used in Step 3 of the SIS/R algorithm above. Alternatively, one can use categorical resampling, which is resampling particles by regarding $\{\widetilde{w}_{t}^{(k)}\}$ as a categorical probability distribution. Many other resampling methods exist (see Douc and Cappé, (2005) for more information).
@@ 5.3 On forecasting for longer horizons
In this section, we briefly discuss what to expect from the forecasting method when the forecasting horizon $h$ becomes longer. Recall that the latent factor series is stationary and follows a stable VAR model. The latent process $\{Z_{t}\}_{t\in\mathbb{Z}}$ is also stationary and its long-term prediction converges to its mean, which is a zero vector. From (B.13) and (B.15), the predicted particles are therefore expected to converge eventually to zero vectors as well. Thus, for longer horizon $h$, we expect
@@
First, in some instances, e.g. Bernoulli and categorical distributions, $\theta_{i}$ represents proportion of count values and is estimated as the corresponding sample proportions. In these cases, for long horizon $h$, we therefore expect our forecast to yield the most likely observed count (modulo the issue of ties). This is the case with the application considered in Section 7. On the other hand, for many other distributions, this observation may not necessarily hold. For example, for the Poisson distribution, the parameter is taken as the sample mean and the most likely count according to this Poisson distribution does not need to be the most likely observed value, let alone be in the sample.
@@
5.4 Model diagnostics
The goodness of fit of a time series model for univariate count data can be assessed through a probability integral transformation (PIT) histogram. See Czado et al., (2009), Kolassa, (2016), Jia et al., (2023). PIT involves computing predictive distributions (see also below). We are not aware of PIT extensions to the multivariate setting, though natural possibilities can certainly be proposed. As with forecasting discussed in Sections 5.15.3, the bigger issue for our model and potentially high-dimensional setting is that obtaining a multivariate predictive distribution would be computationally infeasible. As in those sections, we shall compromise by computing marginal predictive distributions but conditioning on the past of all variables. More specifically, we proceed as described next.
Suppose that our model parameters $\theta,\Lambda,\Psi_{1},\ldots,\Psi_{p},\Sigma_{\varepsilon}$ and $\Sigma_{\eta}$ are estimated as in Sections 3.1, 3.2, and S2. Then, the marginal predictive distribution of $i$th variable is defined as
$P_{i,t}(x_{i})=\mathbb{P}(X_{i,t}\leq x_{i}|X_{1}=x_{1},\ldots,X_{t-1}=x_{t-1})=\mathbb{P}(X_{i,t}\leq x_{i}|x_{0:t-1}).$
It can be approximated similarly to the SMC approximation (5.7) by
$\widehat{P}_{i,t}(x_{i})=\sum_{\ell=0}^{x_{i}}\widehat{\mathbb{E}}_{x_{1:t}}[1_{\{x_{i}=\ell\}}(X_{t})]=\sum_{\ell=0}^{x_{i}}\widehat{\mathbb{E}}_{x_{1:t}}[D_{1_{\{x_{i}=\ell\},t}}(\widehat{Z}_{t|t-1})]=\sum_{\ell=0}^{x_{i}}\sum_{k=1}^{N}\frac{w_{t}^{(k)}}{\Omega_{N,t}}D_{1_{\{x_{i}=\ell\},t}}(\widehat{Z}_{t|t-1}^{(k)}),$
where $w_{t}^{(k)}=w_{t-1}^{(k)}w_{t}(\widehat{Z}_{t|t-1}^{(k)})$ as in (5.6) and $\Omega_{N,t}=\sum_{k=1}^{N}w_{t}^{(k)}$, and by using (5.4),
[…]
$=\Phi\left((\Phi^{-1}(C_{i,\ell})-z)/(\widehat{R}_{t|t-1}^{1/2})_{ii}\right)-\Phi\left((\Phi^{-1}(C_{i,\ell-1})-z)/(\widehat{R}_{t|t-1}^{1/2})_{ii}\right),$
where $\widehat{Z}_{t|t-1}^{(k)}$ is computed as in Section 5.1 using $x_{1},\ldots,x_{t-1}$, and $\widehat{R}_{t|t-1}$ is derived by Kalman recursions in Appendix B.
The (nonrandom) sample PIT for the $i$th component is defined as
[…]
where
$F_{i,t}(u|x_{i,t})=\left\{\begin{array}[]{ll}0,&\textrm{if }u\leq\widehat{P}_{i,t}(x_{i,t}-1),\\ \frac{u-\widehat{P}_{i,t}(x_{i,t}-1)}{\widehat{P}_{i,t}(x_{i,t})-\widehat{P}_{i,t}(x_{i,t}-1)},&\textrm{if }\widehat{P}_{i,t}(x_{i,t}-1)<u<\widehat{P}_{i,t}(x_{i,t}),\\ 1,&\textrm{if }u\geq\widehat{P}_{i,t}(x_{i,t}).\end{array}\right.$
In practice, the sample PIT of the $i$th component is a histogram with the height $\bar{F}_{i}(m/M)-\bar{F}_{i}((m-1)/M)$, $m=1,\ldots,M$, where $M$ is the number of bins (often $M=10$). If the model fits the data well, the sample PIT histogram should be close to a uniform distribution.
@@ 6 Simulation study
In this section, we assess the performance of estimation and forecasting procedures introduced in Sections 3 and 5. We focus on Bernoulli, categorical, Poisson, and negative binomial marginal distributions with several parameter values. Our goal is to check how the performance of our proposed methods compares to that of the traditional PCA methods in the factor modeling literature (e.g., Lam et al., , 2011, Doz et al., , 2011, 2012).
@@ 6.1.1 Estimation for known numbers of factors and lag order
The simulation settings are as follows. First, we fix the order $p=1$ for factor series $\{Y_{t}\}_{t\in\mathbb{Z}}$ in (2.4) and take the number of factor series as $r=2$ or $r=5$. These values are assumed to be given in this section. Then, by following the standard DFM literature (e.g., Doz et al., , 2011), we generate the latent Gaussian DFM (2.3)–(2.4) with parameters:
@@
  • For categorical marginal distributions with $\theta_{i}=(p_{1,i},\ldots,p_{5,i})$ where the number of the categories is given as 5, $\theta_{i}=(0.2,\ldots,0.2)$ (uniform), $(0,0.25,0.5,0.25,0)$ (unimodal), and $(0.45,0,0.1,0,0.45)$ (trimodal) for $i\in I_{1}$, $I_{2}$, and $I_{3}$, respectively.
@@
6.1.2 Selection of the number of factors